Find the general/particular solution of the following Differential Equations.

(Exact D.E)


2.) (2xy-tany)dx+(x²-xSec²y)dy=0


1
Expert's answer
2021-09-07T02:48:55-0400

Solution;

Check for exactness of the equation;

"M=2xy-tany"

"N=x^2-xsec^2y"

"\frac{dM}{dy}=" "2x-sec^2y"

"\frac{dN}{dx}=2x-sec^2y"

Clearly;

"\frac{dM}{dy}=\frac{dN}{dx}"

The equation is exact.

Solution is given as;

"\int_{y=c}Mdx+\int" (Terms of N independent of x)"dy" "=C"

"\int(2xy-tan y)dx+0=C"

"2y\int xdx-tan(y)\int 1dx=C"

"2y×(\frac{x^2}{2})-tany(x)=C"

The general solution is;

"x^2y-xtan(y)=C"